# WAFT: Warping-Alone Field Transforms for Optical Flow

### v1

@@ 5 Experiments
Direct GMFlow [56] 1.74 2.90 9.32 3.80 603 139
CroCoFlow [54] 1.09 2.44 3.64 2.40 57343 6422
DDVM [43] 1.75 2.48
@@
Iterative w/ Cost Volume PWC-Net+ [47] 3.45 4.60 7.72 4.91 101 24
RAFT [50] 1.61 2.86 5.10 3.07 938 141
DIP [66] 1.44 2.83 4.21 2.43 3068 499
GMFlowNet [63] 1.39 2.65 4.79 2.75 1094 244
CRAFT [45] 1.45 2.42 4.79 3.02 2274 483
FlowFormer [13] 1.20 2.12 4.68 2.69 1715 336
GMFlow+ [57] 1.03 2.37 4.49 2.40 1177 250
RPKNet [35] 1.31 2.65 4.64 2.71 137 183
CCMR+ [17] 1.07 2.10 3.86 2.07 12653 999
MatchFlow(G) [6] 1.16 2.37 4.63 2.77 1669 291
Flowformer++[44] 1.07
@@
SEA-RAFT(L) [52] 1.31 2.60 4.30 - 655 108
AnyFlow [21] 1.23 2.44 4.41 2.69 - -
FlowDiffuser [26] 1.02 2.03 4.17 2.82 2466 599
SAMFlow [67]
1.00
2.08 4.49 - 9717 1757
DPFlow [34] 1.04 1.97 3.56 2.12 414 131
Iterative w/ Warping SpyNet [40, 36] 6.64 8.36 35.07 26.71 167 25
FlowNet2 [16] 4.16 5.74 10.41 6.94 230 75
WAFT (Ours)
@@
Table 2: We report endpoint-error (EPE) on Sintel [3] and Fl on KITTI [9]. On KITTI [9], WAFT ranks first on non-occluded pixels and second on all pixels. It also achieves comparable performance to state-of-the-art methods on Sintel (clean). We measure latency on an RTX3090 with batch size 1 and input resolution $540\times 960$.
We report results on Sintel [3], KITTI [9], and Spring [31]. Following existing work [50, 13, 52, 34], for training, we use FlythingChairs [8], FlyingThings [30], HD1K [23], Sintel [3], KITTI [9], Spring [31], and TartanAir [51]. We use the widely adopted metrics: endpoint-error (EPE), 1-pixel outlier rate (1px), percentage of flow outliers (Fl), and weighted area under the curve (WAUC). Definitions can be found in [41, 31, 9, 34].
5.1 Architecture Details
Input Encoder
We use a frozen ViT-S-based DepthAnythingV2 [60] and a ResNet18 [11] in the input encoder. We only use the feature map at the original image resolution from the DepthAnythingV2 head. This DAv2 feature is concatenated with its corresponding image and then fed into ResNet18 for refinement. We upsample the output of ResNet18 to obtain the final image features at 1/2 image resolution using 3 deconv layers.
Recurrent Update Module
We use a modified DPT-Small [39] as the recurrent update module. We concatenate the image features and use a $1\times 1$ conv to obtain the initial hidden state. Since the image features are already 2×2$\times$2 × downsampled, we change the patch size to 8. We set the resolution of the positional embedding to $224\times 224$, and interpolate it for other resolutions. We use $T=5$ iterations in training and inference.
5.2 Training Details
@@
FlowNet2 [16] 6.710 1.040 2.823 90.907
SpyNet [40] 29.963 4.162 12.866 67.150
PWC-Net [46] 82.27 2.288 4.889 45.670
RAFT [50] 6.790 1.476 3.198 90.920
GMA [20] 7.074 0.914 3.079 90.722
FlowFormer [13] 6.510 0.723 2.384 91.679
GMFlow [56] 10.355 0.945 2.952 82.337
RPKNet [35] 4.809 0.657 1.756 92.638
CroCoFlow [54] 4.565 0.498 1.508 93.660
Win-Win [24] 5.371 0.475 1.621 92.270
SEA-RAFT(M) [52] 3.686 0.363 1.347 94.534
DPFlow [34] 3.442 0.340 1.311 94.980
WAFT-540p (Ours)
3.418 0.340 1.280 94.663
WAFT-1080p (Ours)
@@
Table 3: WAFT ranks 1st on Spring [31]. Our 540p model also outperforms existing methods on 1px, EPE, and Fl. denotes the submissions from the Spring team.
@@
Following SEA-RAFT [52], we first pre-train our model on TartanAir [51] for 300k steps, with a batch size of 32 and learning rate $4\times 10^{-4}$. We fine-tune our model on FlyingChairs [62] with the same hyperparameters for 50k steps, and then fine-tune it on FlyingThings [30] with the same hyperparameters for 200k steps. For all submissions, we keep the batch size as 32 by default and reduce the learning rate to $10^{-4}$. For KITTI [9] submission, we fine-tune our model on KITTI(train) [9] for 5k steps. For Sintel [3] submission, we follow previous work to fine-tune our model on the mixture of FlyingThings [30], HD1K [23], KITTI(train) [9], and Sintel(train) [3] for 200k steps. For Spring [31] submissions, we fine-tune two models on Spring(train) [31] for 200k steps: a 1080p model with a batch size of 8 and a 540p model with a batch size of 32.
@@
We first train our model on FlyingChairs [8] for 50k steps, and then fine-tune it on FlyingThings [30] for 50k steps. The batch size is set to 32, and the learning rate is set to $10^{-4}$.
@@ 5.3 Benchmark Results
PWC-Net [46] 2.55 3.93 10.4 33.7
RAFT [50] 1.43 2.71 5.04 17.4
GMA [20] 1.30 2.74 4.69 17.1
SKFlow [48] 1.22 2.46 4.27 15.5
DIP [66] 1.30 2.82 4.29 13.7
EMD-L [5] 0.88 2.55 4.12 13.5
CRAFT [45] 1.27 2.79 4.88 17.5
RPKNet [35] 1.12 2.45 - 13.0
GMFlowNet [63] 1.14 2.71 4.24 15.4
FlowFormer [13] 1.01 2.40 4.09 14.7
Flowformer++ [44] 0.90 2.30 3.93 14.2
CCMR+ [17] 0.98 2.36 - 12.9
MatchFlow(G) [6] 1.03 2.45 4.08 15.6
SEA-RAFT(L) [52] 1.19 4.11 3.62 12.9
AnyFlow [21] 1.10 2.52 3.76 12.4
SAMFlow [67] 0.87
@@
DPFlow [34] 1.02 2.26 3.37 11.1
FlowNet2 [16] 2.02 3.14 10.1 30.4
WAFT (Ours)
@@
Table 4: WAFT achieves the best cross-dataset generalization on KITTI(train), reducing the error by 8%.
Sintel & KITTI
Results are shown in Table 2. On KITTI [9], WAFT achieves the best Fl on non-occluded pixels and the second best on all pixels. WAFT is competitive on Sintel (Final) [3]. On Sintel (Clean), it achieves performance close to state-of-the-art methods. WAFT is up to 4.1×4.1$\times$4.1 × faster than comparable methods, demonstrating its high efficiency.
Spring
Results are shown in Table 3. Benefiting from warping, WAFT can be trained at full resolution (1080p) and achieves the best performance on all 4 metrics. We also train a 540p model following the downsample-upsample protocol of SEA-RAFT [52]; this 540p model outperforms all existing methods on 3 of the 4 metrics: 1px-outlier rate, endpoint-error, and Fl.
Comparison with Existing Warping-based Methods
It appears that warping as an network operation has been largely abandoned by works in the last 8 years and the last time warping-based methods achieved top positions on the leaderboards were around 2017 [40, 16]. WAFT is significant in that it has revisited and revived an idea that has fallen out of favor. Compared to methods that do use warping [40, 16], WAFT reduces endpoint-error (EPE) by at least 59% on Sintel and 67% on Spring [31], while also reducing Fl by at least 67% on KITTI [9] and 56% on Spring [31]. Besides, WAFT reduces 1px-outlier rate by 50% on Spring [31].
5.4 Zero-Shot Evaluation
Following previous work [50, 13, 46], we train our model on FlythingChairs [8] and FlyingThings [30]. Then we evaluate the performance on the training split of Sintel [3] and KITTI [9].
@@ 5.4 Zero-Shot Evaluation
Results are shown in Table 4. WAFT achieves strong cross-dataset generalization. On KITTI(train) [9], WAFT outperforms other methods by a large margin: It improves the endpoint-error (EPE) from 3.37 to 3.10 and Fl from 11.1 to 10.3. On Sintel(train) [3], WAFT achieves performance close to state-of-the-art methods. Compared to previous warping-based method [16], WAFT improves the performance by at least 31%.
@@ 5.5 Ablation Study
Table 5: We report the zero-shot ablation results on Sintel(train) [3] and Spring(sub-val) [31, 52]. The effect of changes can be identified through comparisons with the first row. See Section 5 for details.
We conduct zero-shot ablations in Table 5 on the training split of Sintel [3] and the sub-val split [52] of Spring [31]. In all experiments, the models are trained on FlythingThings [30] for 50k steps, with a batch size of 32 and learning rate $10^{-4}$. The average end-point-error and 1px-outlier rate are reported.
Different Input Encoder
Both pre-trained DepthAnythingV2 [60] feature and ResNet [11] refinement are important to the model performance.
Different Recurrent Update Module
The vision transformer design is crucial to iterative warping. We observe a significant performance drop when replacing the DPT-based recurrent update module with CNNs, highlighting the importance of modeling long-range dependence. This finding may help explain why early deep learning approaches [16, 40] that implemented warping using CNNs underperformed compared to cost-volume-based methods [46, 50, 13].
High-Resolution Indexing
High-resolution indexing into feature maps using current flow estimates remarkably improves performance. We implement variants that index at 1/8 resolution by changing the patch size of DPT to $2\times 2$, and find that high-resolution indexing significantly improves 1px-outlier rate on Spring(sub-val) [31].
We also design a cost-volume-based variant following the common setup [52, 50, 13] with look-up radius $4$ at 1/8 resolution, and find that it performs similarly to the warping counterpart (shown in Table 5) but costs $\times2$ training memory (21.2 GiB vs. 9.5 GiB).
@@
Iterative updates achieve better performance than direct regression within our meta-architecture. We implement direct regression by setting the number of iterations $T=1$. For fair comparison, we scale up the networks to match the computational cost of 5-iteration WAFT. Our results show that WAFT significantly outperforms these direct regression variants, indicating the effectiveness and high efficiency of the iterative paradigm. This finding aligns with the observation that existing direct methods either underperform [56, 53] or require substantially more computational cost [43, 53, 54] compared to iterative approaches [50, 52, 13, 34].
@@
Feature-space warping is more effective than image-space warping, which is commonly used in classic methods [29, 2, 1] and early deep learning methods [16, 40]. Feature-space warping does not need to re-extract features for the warped image in each iteration, significantly saving computational cost while achieving slightly better accuracy.
@@
It is possible to perform iterative updates without warping the features using the current flow estimates. We can simply use the original feature maps as input to the update module. Compared to this baseline, warping has significantly lower error with a negligible cost. This observation aligns with the conclusions of previous work [2], which theoretically justifies the combination of warping and recurrent updates by framing it as a fixed-point iteration algorithm.
@@
Prior work [46, 50, 52, 13] has often used a context encoder that provides an extra input to the update module. Our ablation show that the context encoder is not necessary. The context encoder introduces additional computation overhead, but does not significantly affect performance. Previous work [52] also points out that the context encoder can be regarded as a direct flow regressor, which functions similarly to the first iteration of WAFT.
6 Conclusion
We have introduced Warping-Alone Field Transforms (WAFT), a simple and effective method for optical flow. WAFT is similar to RAFT but replaces cost volume with high-resolution warping, achieving better accuracy and lower memory cost. This design challenges the conventional wisdom that constructing cost volumes is necessary for strong performance.
7 Acknowledgments

### v2

@@ 5 Experiments
We report results on Sintel (Butler et al., 2012), KITTI (Geiger et al., 2013), and Spring (Mehl et al., 2023). Following existing work (Teed & Deng, 2020; Huang et al., 2022; Wang et al., 2024; Morimitsu et al., 2025), for training, we use FlythingChairs (Dosovitskiy et al., 2015), FlyingThings (Mayer et al., 2016), HD1K (Kondermann et al., 2016), Sintel (Butler et al., 2012), KITTI (Geiger et al., 2013), Spring (Mehl et al., 2023), and TartanAir (Wang et al., 2020). We use the widely adopted metrics: endpoint-error (EPE), 1-pixel outlier rate (1px), percentage of flow outliers (Fl), and weighted area under the curve (WAUC). Definitions can be found in (Richter et al., 2017; Mehl et al., 2023; Geiger et al., 2013; Morimitsu et al., 2025).
5.1 Architecture Details
@@
Direct GMFlow (Xu et al., 2022) 1.74 2.90 9.32 3.80 603 139
CroCoFlow (Weinzaepfel et al., 2023) 1.09 2.44 3.64 2.40 57343 6422
DDVM (Saxena et al., 2024) 1.75 2.48
@@
Iterative w/ Cost Volume PWC-Net+ (Sun et al., 2019) 3.45 4.60 7.72 4.91 101 24
RAFT (Teed & Deng, 2020) 1.61 2.86 5.10 3.07 938 141
DIP (Zheng et al., 2022) 1.44 2.83 4.21 2.43 3068 499
GMFlowNet (Zhao et al., 2022) 1.39 2.65 4.79 2.75 1094 244
CRAFT (Sui et al., 2022) 1.45 2.42 4.79 3.02 2274 483
FlowFormer (Huang et al., 2022) 1.20 2.12 4.68 2.69 1715 336
GMFlow+ (Xu et al., 2023b) 1.03 2.37 4.49 2.40 1177 250
RPKNet (Morimitsu et al., 2024) 1.31 2.65 4.64 2.71 137 183
CCMR+ (Jahedi et al., 2024) 1.07 2.10 3.86 2.07 12653 999
MatchFlow(G) (Dong et al., 2023) 1.16 2.37 4.63 2.77 1669 291
Flowformer++(Shi et al., 2023) 1.07
@@
SEA-RAFT(L) (Wang et al., 2024) 1.31 2.60 4.30 - 655 108
AnyFlow (Jung et al., 2023) 1.23 2.44 4.41 2.69 - -
FlowDiffuser (Luo et al., 2024) 1.02 2.03 4.17 2.82 2466 599
SAMFlow (Zhou et al., 2024) 1.00 2.08 4.49 - 9717 1757
DPFlow (Morimitsu et al., 2025) 1.04 1.97 3.56 2.12 414 131
Iterative w/ Warping SpyNet (Ranjan & Black, 2017) 6.64 8.36 35.07 26.71 167 25
FlowNet2 (Ilg et al., 2017) 4.16 5.74 10.41 6.94 230 75
WAFT-DAv2-a1
@@
WAFT-Twins-a2
1.02 2.39 3.53 2.12 1020 290
WAFT-DAv2-a2
0.95
2.33 3.31
2.03
807 240
WAFT-DINOv3-a2
0.94
2.02 3.56 2.13 732 212
Table 2: We report endpoint-error (EPE) on Sintel (Butler et al., 2012), Fl on KITTI (Geiger et al., 2013), and highlight all SOTA performance. On KITTI, WAFT ranks first on non-occluded pixels and second on all pixels. It also achieves state-of-the-art performance on Sintel (clean). We measure the latency on an RTX3090 with batch size 1 and 540p input.
@@
FlowNet2 (Ilg et al., 2017) 6.710 1.040 2.823 90.907
SpyNet (Ranjan & Black, 2017) 29.963 4.162 12.866 67.150
PWC-Net (Sun et al., 2018a) 82.27 2.288 4.889 45.670
RAFT (Teed & Deng, 2020) 6.790 1.476 3.198 90.920
GMA (Jiang et al., 2021) 7.074 0.914 3.079 90.722
FlowFormer (Huang et al., 2022) 6.510 0.723 2.384 91.679
GMFlow (Xu et al., 2022) 10.355 0.945 2.952 82.337
RPKNet (Morimitsu et al., 2024) 4.809 0.657 1.756 92.638
CroCoFlow (Weinzaepfel et al., 2023) 4.565 0.498 1.508 93.660
Win-Win (Leroy et al., 2023) 5.371 0.475 1.621 92.270
SEA-RAFT(M) (Wang et al., 2024) 3.686 0.363 1.347 94.534
DPFlow (Morimitsu et al., 2025) 3.442 0.340 1.311 94.980
WAFT-DAv2-a1-540p
3.418
0.340
1.280
94.663
WAFT-DAv2-a1-1080p
@@
WAFT-Twins-a2
3.268
0.331
1.282
94.786
WAFT-DAv2-a2
3.298
0.304
1.197
94.990
WAFT-DINOv3-a2
3.182
0.325
1.246
95.051
Table 3: WAFT ranks 1st on Spring (Mehl et al., 2023) on all metrics. We highlight all SOTA performance. denotes the submissions from the Spring team.
Input Encoder
We use frozen ImageNet-pretrained Twins-SVT-Large (Chu et al., 2021), depth-pretrained DAv2-S (Yang et al., 2024), and unsupervised-pretrained DINOv3-ViT-S (Siméoni et al., 2025) in input encoders.
Recurrent Update Module
We use a modified DPT-Small (Ranftl et al., 2021) as the recurrent update module. We concatenate the image features and use a $1\times 1$ conv to obtain the initial hidden state. Since the image features are already $2\times$ downsampled, we change the patch size to 8. We set the resolution of the positional embedding to $224\times 224$, and interpolate it for other resolutions. We use $T=5$ iterations in training and inference.
5.2 Training Details
@@
Following SEA-RAFT (Wang et al., 2024), we first pre-train our model on TartanAir (Wang et al., 2020) for 300k steps, with a batch size of 32 and learning rate $4\times 10^{-4}$. We fine-tune our model on FlyingChairs (Zhao et al., 2020a) with the same hyperparameters for 50k steps, and then fine-tune it on FlyingThings (Mayer et al., 2016) for 200k steps. For all submissions, we keep the batch size as 32 by default and reduce the learning rate to $10^{-4}$. For KITTI (Geiger et al., 2013) submission, we fine-tune our model on KITTI(train) for 5k steps. For Sintel (Butler et al., 2012) submission, we follow previous work to fine-tune our model on the mixture of FlyingThings, HD1K (Kondermann et al., 2016), KITTI(train), and Sintel(train) for 200k steps. For Spring (Mehl et al., 2023) submissions, we fine-tune our models on Spring(train) for 200k steps with a batch size of 32. We train an extra 1080p WAFT-DAv2-a1 model with a batch size of 8.
@@
We first train our model on FlyingChairs for 50k steps, and then fine-tune it on FlyingThings for 50k steps. The batch size is set to 32, and the learning rate is set to $10^{-4}$.
@@ 5.3 Benchmark Results
Sintel & KITTI
Results are shown in Table 2. Using a Twins backbone only pre-trained on ImageNet, WAFT ranks second on KITTI and is competitive on Sintel (clean). It outperforms prior cost-volume-based SOTA Flowformer++ (Shi et al., 2023) in both accuracy and efficiency given the same backbone, demonstrating the strength of high-resolution warping. The performance can be further improved with stronger backbones (Yang et al., 2024; Siméoni et al., 2025). Using a depth-pretrained DAv2 (Yang et al., 2024), on KITTI (Geiger et al., 2013), WAFT achieves the best Fl on non-occluded pixels and the second best on all pixels. It also ranks first on Sintel (Clean) (Butler et al., 2012) and competitive on Sintel (Final). WAFT is up to $4.1\times$ faster than comparable methods, demonstrating its high efficiency.
Note that there is an outlier sequence, ‘Ambush 1’, which severely affects the average performance on Sintel (Final) as mentioned in Saxena et al. (2024). We show that WAFT outperforms Flowformer++ on Sintel (Final) with the same Twins backbone when ‘Ambush 1’ is excluded. More details can be found in Table 6.
Spring
Results are shown in Table 3. Following the downsample-upsample protocol (540p) of SEA-RAFT (Wang et al., 2024), WAFT outperforms existing methods on EPE and 1px with a Twins backbone only pre-trained on ImageNet. We also show that WAFT achieves the best performance on all metrics with a depth-pretrained DAv2 backbone. Benefiting from warping, WAFT can be trained at full resolution (1080p) to improve the performance further.
Comparison with Existing Warping-based Methods
It appears that warping as a network operation has been largely abandoned by works in the last 8 years and the last time warping-based methods achieved top positions on the leaderboards were around 2017 (Ranjan & Black, 2017; Ilg et al., 2017). WAFT is significant in that it has revisited and revived an idea that has fallen out of favor. Compared to methods that do use warping (Ranjan & Black, 2017; Ilg et al., 2017), WAFT reduces endpoint-error (EPE) by at least 64% on Sintel and 70% on Spring (Mehl et al., 2023), while also reducing Fl by at least 68% on KITTI (Geiger et al., 2013) and 57% on Spring (Mehl et al., 2023). Besides, WAFT reduces 1px-outlier rate by 52% on Spring (Mehl et al., 2023).
5.4 Zero-Shot Evaluation
@@
PWC-Net (Sun et al., 2018a) 2.55 3.93 10.4 33.7
RAFT (Teed & Deng, 2020) 1.43 2.71 5.04 17.4
GMA (Jiang et al., 2021) 1.30 2.74 4.69 17.1
SKFlow (Sun et al., 2022) 1.22 2.46 4.27 15.5
DIP (Zheng et al., 2022) 1.30 2.82 4.29 13.7
EMD-L (Deng et al., 2023) 0.88 2.55 4.12 13.5
CRAFT (Sui et al., 2022) 1.27 2.79 4.88 17.5
RPKNet (Morimitsu et al., 2024) 1.12 2.45 - 13.0
GMFlowNet (Zhao et al., 2022) 1.14 2.71 4.24 15.4
FlowFormer (Huang et al., 2022) 1.01 2.40 4.09 14.7
Flowformer++ (Shi et al., 2023) 0.90 2.30 3.93 14.2
CCMR+ (Jahedi et al., 2024) 0.98 2.36 - 12.9
MatchFlow(G) (Dong et al., 2023) 1.03 2.45 4.08 15.6
SEA-RAFT(L) (Wang et al., 2024) 1.19 4.11 3.62 12.9
AnyFlow (Jung et al., 2023) 1.10 2.52 3.76 12.4
SAMFlow (Zhou et al., 2024) 0.87
@@
DPFlow (Morimitsu et al., 2025) 1.02 2.26 3.37 11.1
FlowNet2 (Ilg et al., 2017) 2.02 3.14 10.1 30.4
WAFT-DAv2-a1
@@
WAFT-Twins-a2
1.02 2.46
2.98
9.9
WAFT-DAv2-a2
1.01 2.49
3.28
10.9
WAFT-DINOv3-a2
1.28 2.56 3.49 12.9
Table 4: WAFT achieves the best cross-dataset generalization on KITTI(train), reducing the error by 11%. We highlight all SOTA performance.
Following previous work (Teed & Deng, 2020; Huang et al., 2022; Sun et al., 2018a), we train our model on FlyingChairs (Dosovitskiy et al., 2015) and FlyingThings (Mayer et al., 2016). Then we evaluate the performance on the training split of Sintel (Butler et al., 2012) and KITTI (Geiger et al., 2013).
@@
Results are shown in Table 4. WAFT achieves strong cross-dataset generalization. On KITTI (train), WAFT outperforms other methods by a large margin with an ImageNet-pretrained Twins backbone: It improves the endpoint-error (EPE) from 3.37 to 2.98 and Fl from 11.1 to 9.9. On Sintel (train), WAFT achieves performance close to state-of-the-art methods. Compared to the previous warping-based method (Ilg et al., 2017), WAFT improves the performance by at least 31%.
@@ 5.5 Ablation Study
We conduct zero-shot ablations in Table 5 on the training split of Sintel (Butler et al., 2012) and the sub-val split (Wang et al., 2024) of Spring (Mehl et al., 2023) based on WAFT-DAv2-a1. In all experiments, the models are trained on FlythingThings (Mayer et al., 2016) for 50k steps, with a batch size of 32 and learning rate $10^{-4}$. The average EPE and 1px are reported.
Different Input Encoder
Both pre-trained weights and adaptations are important to the performance. Note that the strong performance of WAFT is not merely from advanced backbones. Using a Twins backbone only pre-trained on ImageNet as adopted in Flowformer (Huang et al., 2022), WAFT ranks first on Spring, second on KITTI, and is competitive on Sintel. More details can be found in Table 23, and 4.
Different Recurrent Update Module
The vision transformer design is crucial to iterative warping. We observe a significant performance drop when replacing the DPT-based recurrent update module with CNNs, highlighting the importance of modeling long-range dependence. This finding may help explain why early deep learning approaches (Ilg et al., 2017; Ranjan & Black, 2017) that implemented warping using CNNs underperformed compared to cost-volume-based methods (Sun et al., 2018a; Teed & Deng, 2020; Huang et al., 2022).
High-Resolution Indexing
High-resolution indexing into feature maps using current flow estimates remarkably improves performance. We implement variants that index at 1/8 resolution by changing the patch size of DPT to $2\times 2$, and find that high-resolution indexing significantly improves 1px-outlier rate on Spring(sub-val) (Mehl et al., 2023).
We also design a cost-volume-based variant following the common setup (Wang et al., 2024; Teed & Deng, 2020; Huang et al., 2022) with look-up radius $4$ at 1/8 resolution, and find that it performs similarly to the warping counterpart (shown in Table 5) but costs $2.2\times$ training memory (21.2 GiB vs. 9.5 GiB).
@@
Table 5: We report the zero-shot ablation results on Sintel(train) (Butler et al., 2012) and Spring(sub-val) (Mehl et al., 2023; Wang et al., 2024). The effect of changes can be identified through comparisons with the first row. See Section 5.5 for details.
@@
Iterative updates achieve better performance than direct regression within our meta-architecture. We implement direct regression by setting the number of iterations $T=1$. For fair comparison, we scale up the networks to match the computational cost of 5-iteration WAFT. Our results show that WAFT significantly outperforms these direct regression variants, indicating the effectiveness and high efficiency of the iterative paradigm. This finding aligns with the observation that existing direct methods either underperform (Xu et al., 2022; Weinzaepfel et al., 2022) or require substantially more computational cost (Saxena et al., 2024; Weinzaepfel et al., 2022; 2023) compared to iterative approaches (Teed & Deng, 2020; Wang et al., 2024; Huang et al., 2022; Morimitsu et al., 2025).
@@
Feature-space warping is more effective than image-space warping, which is commonly used in classic methods (Ma et al., 2022; Brox et al., 2004; Black & Anandan, 1996) and early deep learning methods (Ilg et al., 2017; Ranjan & Black, 2017). Feature-space warping does not need to re-extract features for the warped image in each iteration, significantly saving computational cost while achieving slightly better accuracy.
@@
It is possible to perform iterative updates without warping the features using the current flow estimates. We can simply use the original feature maps as input to the update module. Compared to this baseline, warping has significantly lower error with a negligible cost. This observation aligns with the conclusions of previous work (Brox et al., 2004), which theoretically justifies the combination of warping and recurrent updates by framing it as a fixed-point iteration algorithm.
@@
Prior work (Sun et al., 2018a; Teed & Deng, 2020; Wang et al., 2024; Huang et al., 2022) has often used a context encoder that provides an extra input to the update module. Our ablation show that the context encoder is not necessary. The context encoder introduces additional computation overhead, but does not significantly affect performance. Previous work (Wang et al., 2024) also points out that the context encoder can be regarded as a direct flow regressor, which functions similarly to the first iteration of WAFT.
Acknowledgments